# SwiftBrush[Uncaptioned image]: One-Step Text-to-Image Diffusion Model with Variational Score Distillation

### v1

@@ 1 Introduction
Several approaches have been proposed to improve the inference speed of diffusion-based text-to-image generation. However, the most effective direction is time-step distillation, which aims to reduce the number of sampling steps of the diffusion model with the network backbone unchanged and minimum reduction in generation quality. Starting from the pioneering work [35], more advanced and effective distillation methods have been proposed, and a few recent ones could train efficient student networks that only require one forward step to synthesize images from input captions. In this work, we are interested in the distillation-based approach for one-step text-to-image generation. Particularly, Meng et al. [23] employ a two-stage distillation, where the student tries to first match its output with the classifier-free guided output of the teacher and then bootstrap itself to reduce the inference steps gradually. Such a method is complex and time-consuming, requiring at least $108$ A100 GPU days [19]. Based on the work of Song et al. [40], Latent Consistency Model (LCM) [22] enforces the consistency of all points on the ODE trajectory in the latent space of Stable Diffusion [41] and unlike [23], they directly combine the two distillation stages into one. Although both Guided Distillation and LCM significantly reduce the number of inference steps to $2$ or $4$, their 1-step inference yields blurry and unsatisfying results. More sophisticated, Instaflow [19] proposes a technique called “reflow” to straighten the sampling trajectory of the teacher before distilling. While their proposed techniques proven to be superior to the original diffusion framework in one-step text-to-image distillation, InstaFlow requires an expensive 4-stage training schedule. Also, it is noteworthy that the effectiveness of these methods relies heavily on a large number of text-image pairs, which is not always the case due to the limited accessibility of such data.
@@
As a result, our method has successfully enabled the generation of high-quality images with remarkable details. Quantitatively, our one-step model achieves promising zero-shot results, 16.67 for the FID score and 0.29 for the CLIP score on the MS COCO 2014 dataset (30K images). Notably, this is the first time a distilled one-step model outperforms the previous approaches without a single image used for training. A demonstration of generated images from our distilled model is shown in LABEL:fig:teaser.
@@ 2 Related Work
Text-to-3D Generation. Instead of learning a generative model directly from 3D supervision, many research works have utilized the rich knowledge of 2D prior for 3D generation tasks. For instance, one of the earliest works, Dream Fields [12], leverages CLIP [27] to guide the generated images so that rendered images from many camera views highly align with the text caption. On the other hand, pioneer works like DreamFusion [26] and Score Jacobian Chaining [42] (SJC) propose two different but equivalent frameworks where a 2D text-to-image generative teacher is used to generate 3D objects from textual descriptions. Subsequent works seek to improve these text-to-3D methods in various ways. Among them, Magic3D [17] and ProlificDreamer [43] are two remarkable works that effectively enhance the generation quality of 3D assets. Magic3D consists of two stages where they first obtain a low-resolution 3D representation via a DreamFusion-style optimization and then refine its texture to get the final high-fidelity 3D mesh. Meanwhile, ProlificDreamer proposes variational score distillation involving a secondary teacher to bridge the gap between the teacher and the 3D NeRF. Influenced by the latter work, our method adapts the same technique to develop a one-step text-to-image generator.
Knowledge Distillation introduced by Hinton et al. [8] is one of the methods falling under the umbrella of transfer learning. This algorithm draws its main inspiration from the human learning process, where knowledge is conveyed from a more knowledgeable teacher to a less knowledgeable student. In the context of diffusion models, works have been proposed for faster learning of student models from a pre-trained diffusion teacher. The most straightforward approach involves direct distillation [21], wherein a student model is supervised from the teacher’s output after sampling for 50-1000 steps, which can be prohibitive. Recent work [40, 35] avoids the lengthy sampling during distillation by various bootstrapping methods. Afterward, a series of methods [22, 19, 23] adapt these methods into text-to-image settings and inherit their limitation of being image-dependent in training. In contrast, a concurrent work [7] learns a time-conditioned model to predict the outcome of the teacher at any given time step, all without the need for image supervision. Likewise, our approach is similarly image-free yet achieves significantly better results and offers a much simpler distillation design.
@@ 3.1 Preliminary
Conversely, the backward process starts from $x_{T}\sim\mathcal{N}(0,I)$ and iteratively denoises for $T$ time steps with each step $t$ converts from noisy variable $x_{t+1}$ to less-noisy variable $x_{t}$, guided by a model $\epsilon_{\psi}$ that predicts the noise $\epsilon$ in Eq. 1. The network weights $\psi$ are trained by minimizing:
@@
While common diffusion models operate in the pixel space, Latent Diffusion Models (LDMs)[32] model the diffusion process in the latent space generated by a pre-trained regularized autoencoder, which typically features representations of smaller resolution, thus enhancing both training and inference efficiency.
@@
$\mathcal{L}_{diff}(\psi)=\mathbb{E}_{t,\color[rgb]{0,1,1}y\color[rgb]{0,0,0},\epsilon\in\mathcal{N}(0,1)}\|\epsilon_{\psi}(x_{t},t,\color[rgb]{0,1,1}y\color[rgb]{0,0,0})-\epsilon\|_{2}^{2}$ (3)
@@
Score Distillation Sampling (SDS) is a distillation technique for pre-trained diffusion models, effectively applied in generating 3D assets [26, 42, 17]. It utilizes a pre-trained text-to-image diffusion model, which predicts diffusion noise from text condition $y$, denoted as $\epsilon_{\psi}(x_{t},t,y)$. The method optimizes a single 3D NeRF, parameterized by $\theta$, to align with a given text prompt. Given camera parameters $c$, a differentiable rendering function $g(\cdot,c)$ is used to render an image at the camera view $c$ from the 3D NeRF. Here, the rendered image $g(\theta,c)$ is utilized to optimize the weight $\theta$ through a loss function whose gradient can be approximated by:
@@
$\begin{split}\nabla_{\theta}\mathcal{L}_{VSD}=\mathbb{E}_{t,\epsilon,c}[&w(t)(\epsilon_{\psi}(x_{t},t,y)\\ &-\color[rgb]{0,1,1}\epsilon_{\phi}(x_{t},t,y,c)\color[rgb]{0,0,0})\frac{\partial g(\theta,c)}{\partial\theta}]\end{split}$ (5)
@@ 3.2 SwiftBrush
Motivation. While SDS and VSD is designed explicitly for the text-to-3D generation task, they loosely connect to that goal via the rendered image $g(\theta,c)$ of the 3D NeRF. As a matter of fact, instead of using NeRF rendering, we can replace it with any function that output 2D images to suit our needs. Inspired by this motivation, we propose to substitute the NeRF rendering $g(\theta,c)$ with a text-to-image generator that can directly synthesize a text-guided image in one step, effectively converting the text-to-3D generation training into one-step diffusion model distillation.
Design Space. We adopt the same approach as [43], with modifications to the design space to suit our task better. To begin with, we make use of two teacher models: a pretrained text-to-image teacher $\epsilon_{\psi}$ and one additional LoRA teacher $\epsilon_{\phi}$. Also, we remove the conditioning on camera view $c$ from the LoRA teacher as it is unnecessary in our case, and we use classifier-free guidance for both teachers. Then, we replace the NeRF, which is overfitted for a specific user-provided prompt in a text-to-3D setting, with a generalized one-step text-to-image student model $f_{\theta}(z,y)$. Our student model $f_{\theta}$ takes a random Gaussian noise $z$ and a text prompt $y$ as inputs. Both the LoRA teacher and student model are initialized with the weight of the text-to-image teacher. Next, we train both the student model and the LoRA teacher alternately using Eq. 5 and Eq. 6 while freezing the text-to-image teacher. A pseudo-code and system figure can be viewed in Alg. 1 and Fig. 1.
Student Parameterization. Given a pretrained text-to-image diffusion model $\epsilon_{\theta}$, it is possible to directly use its output for the student model, i.e., $f_{\theta}(z,y)=\epsilon_{\theta}(z,T,y)$, where $T$ is the maximum timestep of the pretrained model. However, in our case, the pretrained model of choice is Stable Diffusion, which is inherently designed to predict the added noise $\epsilon$. In contrast, our objective is to refine the student model so that it predicts a clean, noise-free image $x_{0}$. Therefore, there is a large domain gap between what we want the student to learn and the output of the student. For the sake of easy training, we empirically re-parameterize the student output as follows:
@@
Figure 1: SwiftBrush overview. Our system allows training a one-step text-to-image student network from a frozen pretrained teacher with the help of an additional trainable LORA teacher. The student network inputs a text prompt and random noise to synthesize an image. The image is then added noise and sent to two teachers, alongside the prompt and a randomly drawn timestep, to compute the gradient of VSD loss, which backpropagates back to the student. Apart from that, the LoRA teacher is also updated using diffusion loss. Similar to ProlificDreamer [43], we update the LoRA teacher (green flow) and the student (red flow) alternatively .
@@ 4.1 Experimental Setup
Experiment Protocols. In these experiments, we rely on Stable Diffusion v2.1 [41] as our teacher model, and whenever possible, we directly utilize the reported values from other research. As for InstaFlow and LCM, we use the their provided pretrained models to generate images for the quantitative evaluation. Conversely, we are compelled to re-implement BOOT since they did not provide any metrics for their text-to-image models, and their codebase remains entirely closed.
Implementation Details. Similar to prior research in the literature, we use student models with the same architectures as those of the teachers and initialize all of the student parameters with the teacher model. In the training process, specific configuration parameters are used. For the LoRA teacher model, we apply a learning rate of 1e-3, a LoRA rank of 64, and an alpha value of 108. Meanwhile, the student model’s learning rate is set at 1e-6 with exponential moving average (EMA) enabled. Adam optimizer [14] and a batch size of 64 are used to train both the student and LoRA teacher for around 65K iterations. Thanks to the image-free nature of SwiftBrush, we do not require access to any image training set, and only the prompts are needed to distill the generator. Here, we utilize 4M captions from the large-scale text-image dataset JourneyDB [25]. Regarding BOOT, we closely follow their implementation guidelines to the best of our abilities and train the student model until its convergence, except that we use Stable Diffusion v2.1 instead of DeepFloyd IF [5] as the teacher.
@@ 4.2 Results
Quantitative Results. We first evaluate the proposed method on standard text-to-image generation benchmarks. The quantitative comparison with the previous distillation, both image-dependent and image-free, and the original Stable Difusion teacher is shown in Tab. 1. Despite being trained with only text captions, SwiftBrush significantly improves the quality of 1-step inference and achieves better photo-realism as well as text-image alignment against other methods, with exceptions of InstaFlow and Stable Diffusion (25 sampling steps) in FID score. However, it is worth noting that Instaflow is so complex, requiring a 4-stage distillation process, while ours is end-to-end and straightforward. We further conduct a quantitative evaluation with the HPsv2 score, a metric trained to capture human preference. Again, we obtain the best score in all aspects among all mentioned works, which is only a slight degradation compared to the original Stable Diffusion teacher (Tab. 2). Moreover, according to the anonymous survey carried out by us (Fig. 2), our method consistently outperforms Instaflow in terms of image fidelity and text-to-image accuracy.
@@
Qualitative Results. We report the visual comparison between our method and the mentioned baseline methods for the one-step text-to-image generation task. As seen in Fig. 3, LCM outputs blurry and over-smoothed images in the 1-step regime. BOOT faithfully produces images that match well with the text caption, but its results still fall short in terms of fidelity. In contrast, a much better quality can be observed in the image produced by Instaflow and SwiftBrush. InstaFlow, nevertheless, cannot handle complex or unusual input prompts properly, which can be noticed in the third and fifth images at the third column of Fig. 3. On top of that, it occasionally produces non-realistic results even for simple prompts like in Fig. 3 at the fourth image of the third column. When, in fact, the outputs generated by SwiftBrush are not only of exceptional quality but also exhibit a remarkable alignment between text descriptions and the corresponding images. Moreover, our results are comparable with those sampled from Stable Diffusion with 25 steps, whereas it is understandable that with only one step, Stable Diffusion cannot create outputs of good quality.
Figure 3: Sample images generated by LCM, BOOT, Instaflow, Ours, SD with 25 sampling steps ($*{}^{*}$) and SD with 1 sampling step (${}^{\dagger}$). Images at the same row are sampled from the same text prompt, while images at the same column are from the same model. We use DPM-Solver [20] to sample Stable Diffusion
@@
Guided ${}^{\dagger}$ 1 37.3 0.27
${}^{\dagger}$ 1 35.56 0.24
[…]
${}^{\ddagger}$ 1 48.20 0.26
@@
SD $*{}^{*}$ 25 13.45 0.23
SD $*{}^{*}$ 1 202.14 0.06
[…]
@@
SD $*{}^{*}$ 27.48 26.89 26.86 27.46
Table 2: Comparison of our method against other works based on HPSv2 score in 1-step regime. ${}^{\dagger}$ means that we obtain the score using the provided pretrained models of the corresponding papers. ${}^{\ddagger}$ means that we re-implement the work and report the score by ourselves. $*{}^{*}$ means that we obtain the score from [45]. Bold and underlined numbers are best and second best, respectively.

### v4

@@ 1 Introduction
Several approaches have been proposed to improve the inference speed of diffusion-based text-to-image generation. However, the most effective direction is time-step distillation, which aims to reduce the number of sampling steps of the diffusion model with the network backbone unchanged and minimum reduction in generation quality. Starting from the pioneering work [37], more advanced and effective distillation methods have been proposed, and a few recent ones could train efficient student networks that only require one forward step to synthesize images from input captions. In this work, we are interested in the distillation-based approach for one-step text-to-image generation. Particularly, Meng et al. [25] employ a two-stage distillation , where the student tries to first match its output with the classifier-free guided output of the teacher and then bootstrap itself to reduce the inference steps gradually. Such a method is complex and time-consuming, requiring at least $108$ A100 GPU days [20]. Based on the work of Song et al. [42], Latent Consistency Model (LCM) [23] enforces the consistency of all points on the ODE trajectory in the latent space of Stable Diffusion (SD) [43] and unlike [25], they directly combine the two distillation stages into one. Although both Guided Distillation and LCM significantly reduce the number of inference steps to $2$ or $4$, their 1-step inference yields blurry and unsatisfying results. More sophisticated, Instaflow [20] proposes a technique called “reflow” to straighten the sampling trajectory of the teacher before distilling. While their proposed techniques proven to be superior to the original diffusion framework in one-step text-to-image distillation, InstaFlow requires an expensive 4-stage training schedule. Also, it is noteworthy that the effectiveness of these methods relies heavily on a large number of text-image pairs, which is not always the case due to the limited accessibility of such data.
@@
As a result, our method has successfully enabled the generation of high-quality images with remarkable details. Quantitatively, our one-step model achieves promising zero-shot results, 16.6 for the FID score and 0.29 for the CLIP score on the MS COCO 2014 dataset (30K images). Notably, this is the first time a distilled one-step model outperforms the previous approaches without a single image used for training. A demonstration of generated images from our distilled model is shown in LABEL:fig:teaser.
@@ 2 Related Work
Text-to-3D Generation. Instead of learning a generative model directly from 3D supervision, many research works have utilized the rich knowledge of 2D prior for 3D generation tasks. For instance, one of the earliest works, Dream Fields [13], leverages CLIP [29] to guide the generated images so that rendered images from many camera views highly align with the text caption. On the other hand, pioneer works like DreamFusion [28] and Score Jacobian Chaining [44] (SJC) propose two different but equivalent frameworks where a 2D text-to-image generative prior is used to generate 3D objects from textual descriptions. Subsequent works seek to improve these text-to-3D methods in various ways. Among them, Magic3D [18] and ProlificDreamer [45] are two remarkable works that effectively enhance the generation quality of 3D assets. Magic3D consists of two stages where they first obtain a low-resolution 3D representation via a DreamFusion-style optimization and then refine its texture to get the final high-fidelity 3D mesh. Meanwhile, ProlificDreamer proposes Variational Score Distillation (VSD) involving a secondary teacher to bridge the gap between the teacher and the 3D NeRF. Influenced by the latter work, our method adapts the same technique for one-step text-to-image generation. Similarly, Score GAN [6] and Diff-Instruct [24] use a VSD-like loss to distill a pretrained diffusion model to a one-step generator.
Knowledge Distillation introduced by Hinton et al. [9] is one of the methods falling under the umbrella of transfer learning. This algorithm draws its main inspiration from the human learning process, where knowledge is conveyed from a more knowledgeable teacher to a less knowledgeable student. In the context of diffusion models, works have been proposed for faster sampling of student models from a pre-trained diffusion teacher. The most straightforward approach involves direct distillation [22], wherein a student model is supervised from the teacher’s output after sampling for 50-1000 steps, which can be prohibitive. Recent work [42, 37] avoids the lengthy sampling during distillation by various bootstrapping methods. Afterward, a series of methods [23, 20, 25] adapt these methods into text-to-image settings and inherit their limitation of being image-dependent in training. In contrast, a concurrent work [8] learns a time-conditioned model to predict the outcome of the teacher at any given time step, all without the need for image supervision. Likewise, our approach is similarly image-free yet achieves significantly better results and offers a much simpler distillation design.
@@ 3.1 Preliminary
Conversely, the backward process starts from $x_{T}\sim\mathcal{N}(0,I)$ and iteratively denoises for $T$ steps where at each step $t$, noisy variable $x_{t+1}$ is converted to less-noisy variable $x_{t}$, guided by a model $\epsilon_{\psi}$ that predicts the noise $\epsilon$ in Eq. 1. The network weights $\psi$ are trained by minimizing:
@@
While common diffusion models operate in the pixel space, Latent Diffusion Models (LDMs)[34] model the diffusion process in the latent space of a pre-trained regularized autoencoder, which features representations of smaller size, thus enhancing both training and inference efficiency.
@@
$\mathcal{L}_{diff}(\psi)=\mathbb{E}_{t,y,\epsilon\in\mathcal{N}(0,1)}\|\epsilon_{\psi}(x_{t},t,y)-\epsilon\|_{2}^{2}$ (3)
@@
Score Distillation Sampling (SDS) is a distillation technique effectively applied in generating 3D assets [28, 44, 18]. It utilizes a pre-trained text-to-image diffusion model, which predicts diffusion noise from text condition $y$, denoted as $\epsilon_{\psi}(x_{t},t,y)$. The method optimizes a single 3D NeRF, parameterized by $\theta$, to align with a given text prompt. Given camera parameters $c$, a differentiable rendering function $g(\cdot,c)$ is used to render an image at the camera view $c$ from the 3D NeRF. Here, the rendered image $g(\theta,c)$ is utilized to optimize the weight $\theta$ through a loss function whose gradient can be approximated by:
@@
$\begin{split}\nabla_{\theta}\mathcal{L}_{VSD}=\mathbb{E}_{t,\epsilon,c}[&w(t)(\epsilon_{\psi}(x_{t},t,y)\\ &-\epsilon_{\phi}(x_{t},t,y,c))\frac{\partial g(\theta,c)}{\partial\theta}]\end{split}$ (5)
@@ 3.2 SwiftBrush
Motivation. While SDS and VSD are designed explicitly for the text-to-3D generation task, they loosely connect to that goal via the rendered image $g(\theta,c)$ of the 3D NeRF. As a matter of fact, instead of using NeRF rendering, we can replace it with any function that outputs 2D images to suit our needs. Inspired by this motivation, we propose to substitute the NeRF rendering $g(\theta,c)$ with a text-to-image generator that can directly synthesize a text-guided image in one step, effectively converting the text-to-3D generation training into one-step diffusion model distillation.
Design Space. We adopt the same approach as ProlificDreamer [45], with modifications to the design space to suit our task better. To begin with, we make use of two teacher models: a pretrained text-to-image teacher $\epsilon_{\psi}$ and one additional LoRA teacher $\epsilon_{\phi}$. Also, we remove the conditioning on camera view $c$ from the LoRA teacher as it is unnecessary in our case, and we use classifier-free guidance for both teachers. Then, we replace the NeRF, which is overfitted for a specific user-provided prompt in a text-to-3D setting, with a generalized one-step text-to-image student model $f_{\theta}(z,y)$. Our student model $f_{\theta}$ takes a random Gaussian noise $z$ and a text prompt $y$ as inputs. Both the LoRA teacher and student model are initialized with the weight of the text-to-image teacher. Next, we train both the student model and the LoRA teacher alternately using Eq. 5 and Eq. 6 while freezing the text-to-image teacher. A pseudo-code and system figure can be viewed in Alg. 1 and Fig. 1.
Student Parameterization. Given a pretrained text-to-image diffusion model $\epsilon_{\theta}$, it is possible to directly use its output for the student model, i.e., $f_{\theta}(z,y)=\epsilon_{\theta}(z,T,y)$, where $T$ is the maximum timestep of the pretrained model. However, in our case, the pretrained model of choice is Stable Diffusion, which is inherently designed to predict the added noise $\epsilon$. In contrast, our objective is to refine the student model so that it predicts a clean and noise-free $x_{0}$. Therefore, such a naive method results in a large domain gap between what we want the student to learn and the output of the student. For the sake of easy training, we empirically re-parameterize the student output as follows:
@@
Figure 1: SwiftBrush overview. Our system allows training a one-step text-to-image student network from a frozen pretrained teacher with the help of an additional trainable LORA teacher. The student network takes inputs a text prompt and random noise. The output of the student is then added noise and sent to two teachers, alongside the prompt and a randomly drawn timestep, to compute the gradient of VSD loss, which backpropagates back to the student. Apart from that, the LoRA teacher is also updated using diffusion loss. Similar to ProlificDreamer [45], we update the LoRA teacher (green flow) and the student (red flow) alternatively.
@@ 4.1 Experimental Setup
Experiment Protocols. In these experiments, we rely on Stable Diffusion v2.1 [43] for our two teacher models, and whenever possible, we directly utilize the reported values from other research. As for InstaFlow and LCM, we use the their provided pretrained models to generate images for evaluation. Conversely, we are compelled to re-implement BOOT since they did not provide any metrics for their text-to-image models, and their codebase remains entirely closed.
Implementation Details. Similar to prior research, we use student models with the same architectures as those of the teachers and initialize all of the student parameters with the teacher model. During training, a guidance scale of $4.5$ is used for both teachers. For the LoRA teacher model, we apply a learning rate of 1e-3, a LoRA rank of 64, and an alpha value of 108. Meanwhile, the student model’s learning rate is set at 1e-6 with exponential moving average (EMA) used every iteration. Adam optimizer [15] and a batch size of 64 (16 per GPU) are used to train both the student and LoRA teacher for around 65K iterations. Thanks to the image-free nature of SwiftBrush, we do not require access to any image training set, and only the prompts are needed to distill the generator. Here, we utilize 4M captions from the large-scale text-image dataset JourneyDB [27], which we subsequently deduplicated to 2M. Regarding BOOT, we closely follow their implementation guidelines to the best of our abilities and train the student model until its convergence, except that we use Stable Diffusion v2.1 instead of DeepFloyd IF [5] as the teacher. Here, its blurry results (Fig. 3) compared to the original work could be due to the different teacher.
@@ 4.2 Results
Quantitative Results. We first evaluate the proposed method on standard text-to-image generation benchmarks. The quantitative comparison with the previous distillation, both image-dependent and image-free, and the original Stable Diffusion teacher is shown in Tab. 1. Despite being trained with only text captions, SwiftBrush significantly improves the quality of 1-step inference and achieves better photo-realism as well as text-image alignment against other methods, with exceptions of InstaFlow and Stable Diffusion (25 sampling steps) in FID score. However, it is worth noting that Instaflow is so complex, requiring a 4-stage distillation process, while ours is end-to-end and straightforward. We further conduct a quantitative evaluation with the HPSv2 score, a metric trained to capture human preference. Again, we obtain the best score in all aspects among all mentioned works, which is only a slight degradation compared to the original Stable Diffusion teacher (Tab. 2). Moreover, according to the anonymous survey carried out by us (Fig. 2), our method outperforms Instaflow in terms of image fidelity and text-to-image accuracy.
@@
Qualitative Results. We report the visual comparison between our method and the mentioned baseline methods for the one-step text-to-image generation task. As seen in Fig. 3, LCM outputs blurry and over-smoothed images in the 1-step regime. BOOT faithfully produces images that match well with the text caption, but its results still fall short in terms of fidelity. In contrast, a much better quality can be observed in the image produced by Instaflow and SwiftBrush. InstaFlow, nevertheless, cannot handle complex or unusual input prompts properly, which can be noticed in the third and fifth images at the third column of Fig. 3. On top of that, it occasionally produces non-realistic results even for simple prompts like in Fig. 3 at the fourth image of the third column. When, in fact, the outputs generated by SwiftBrush are not only of exceptional quality but also exhibit a remarkable alignment between text descriptions and the corresponding images. Moreover, our results are comparable with those sampled from Stable Diffusion with 25 steps; whereas it is understandable that with only one step, Stable Diffusion cannot create outputs of good quality.
Figure 3: Sample images generated by LCM, BOOT, Instaflow, Ours, SD with 25 sampling steps (SD) and SD with 1 sampling step (SD). Images at the same row are sampled from the same text prompt, while images at the same column are from the same model. We use DPM-Solver [21] with guidance scale $7.5$ to sample for Stable Diffusion
@@
Guided Distillation 1 37.3 0.27
LCM 1 35.56 0.24
Instaflow 1 13.10 0.28§
BOOT 1 48.20 0.26
@@
SD 2.1 25 13.45 0.30
SD 2.1 1 202.14 0.08
Table 1: Comparison of our method against other works based on FID metric and CLIP score on the COCO 2014 dataset. means that we obtain the numbers from the corresponding papers. § means that we obtain the numbers using the provided pretrained models of the corresponding papers. means that we re-implement the work and report the numbers. means that we report the number using the pretrained models. We use DPM-Solver [21] with guidance scale $7.5$ to sample for SD. Bold and underlined numbers are best and second best for one-step models.
@@
SD 2.1 27.48 26.89 26.86 27.46
Table 2: Comparison of our method against other works based on HPSv2 score in 1-step regime. means that we obtain the score using the provided pretrained models of the corresponding papers. means that we re-implement the work and report the score by ourselves. means that we obtain the score from [47]. Bold and underlined numbers are best and second best, respectively.